SourMunchkin7806:

Find the cube roots of 27(cos 279° + i sin 279°).

SourMunchkin7806:

@Hero

SourMunchkin7806:

@Vocaloid

Vocaloid:

r^(1/n) * [cos(0 + 2pi *k)/n + i * sin(0 + 2pi*k)/n) where n = 3 (since we have a cube root), r = 27, and k = 0, 1, and 2 you must plug in k = 0, 1, and 2 separately to generate three roots

SourMunchkin7806:

im supposed to show how i get the answer though like step by step can you help me with that

SourMunchkin7806:

like i know the answer i just dont know how to explain it all

SourMunchkin7806:

the answer i get is 3(cos 93 + i sin 93)

Vocaloid:

well first we start w/ the formula r^(1/n) * [cos(0 + 2pi *k)/n + i * sin(0 + 2pi*k)/n) i don't think they'd want you to derive this formula, it's pretty complex plugging in r and n gives us 27^(1/3) * [cos(0 + 2pi *k)/3 + i * sin(0 + 2pi*k)/3) then first root: k = 0 gives us 27^(1/3) * [cos(0 + 2pi *0)/3 + i * sin(0 + 2pi*0)/3) second root k = 1 gives us 27^(1/3) * [cos(0 + 2pi *1)/3 + i * sin(0 + 2pi*1)/3) third root k = 2 gives us 27^(1/3) * [cos(0 + 2pi *2)/3 + i * sin(0 + 2pi*2)/3) showing you plugging in the numbers and doing any algebraic simplification should be sufficient

SourMunchkin7806:

So what do i put just that formula with me explaining what i plugged in?

Vocaloid:

yes

SourMunchkin7806:

Thank you!

SourMunchkin7806:

So our formula we use is, r^(1/n) * [cos(0 + 2pi *k)/n + i * sin(0 + 2pi*k)/n) So first we plug in r and n gives which gives us this, 27^(1/3) * [cos(0 + 2pi *k)/3 + i * sin(0 + 2pi*k)/3) So now we plug in all of our roots (0,1,2) k = 0 gives us 27^(1/3) * [cos(0 + 2pi *0)/3 + i * sin(0 + 2pi*0)/3) k = 1 gives us 27^(1/3) * [cos(0 + 2pi *1)/3 + i * sin(0 + 2pi*1)/3) k = 2 gives us 27^(1/3) * [cos(0 + 2pi *2)/3 + i * sin(0 + 2pi*2)/3) So our cube roots for this equation are (0,1, and 2)

SourMunchkin7806:

does that look good?

Vocaloid:

yes they probably expect you to simplify the angles, too ex: cos(0 + 2pi *0)/3 becomes cos (0)

Vocaloid:

hold on

SourMunchkin7806:

ok

Vocaloid:

ah, I forgot the theta value they give you is 279 not 0 then first root: k = 0 gives us 27^(1/3) * [cos(279 + 2pi *0)/3 + i * sin(0 + 2pi*0)/3) second root k = 1 gives us 27^(1/3) * [cos(279 + 2pi *1)/3 + i * sin(0 + 2pi*1)/3) third root k = 2 gives us 27^(1/3) * [cos(279 + 2pi *2)/3 + i * sin(0 + 2pi*2)/3)

SourMunchkin7806:

ok sweet thanks